# A Smooth Approximation Framework for Weakly Convex Optimization

### v1

@@ 5.2 Smooth minimization with inexact proximal point
We see from Section 5.1 that directly applying smoothing does not surpass the complexity of the subgradient method [davis2019stochastic]. The main issue is that the simple gradient descent algorithm is ineffective in dealing with the ill-conditioned smoothed problem arising from the asymmetry between the lower curvature $\bar{\rho}$ and the upper curvature $L_{\eta}$. To further improve the convergence rate, we propose solving the smooth problem using the proximal point method (Algorithm 2), which turns the original problem into a sequence of strongly convex subproblems.
@@ 2   Compute an approximate solution $\mathbf{x}^{k}$ to the following problem:
The subproblems (36) can be solved more efficiently with acceleration. For convenience, we denote
@@
Finally, let ${\mathbf{x}}^{\star}_{\eta}$ be an optimal solution to the smoothed problem. We assume that the proximal subproblem (36) is inexactly solved: the suboptimality of $\mathbf{x}^{k}$ consists of a relative error to the initial suboptimality plus an absolute error term:
@@
$\min_{1\leq k\leq K}\mathbb{E}[\lVert\nabla\phi_{\eta}^{\hat{\rho}}(\mathbf{x}^{k})\rVert^{2}]\leq\frac{2\hat{\rho}^{2}}{\hat{\rho}-\bar{\rho}}\frac{(1+\lambda)\big[\phi(\mathbf{x}^{0})-\phi(\mathbf{x}^{\star})+R\eta\big]+\sum_{k=1}^{K-1}\mathbb{E}[\zeta_{k}]}{(1-\lambda)K}.$
@@
$(1-\lambda)\sum_{k=2}^{K+1}\mathbb{E}\big[{\hat{\phi}}_{k}(\mathbf{x}^{k-1})-{\hat{\phi}}_{k}(\hat{\mathbf{x}}^{k})\big]$
@@
$\leq(1+\lambda)\big[{\phi_{\eta}}(\mathbf{x}^{0})-{\phi_{\eta}}({\mathbf{x}}^{\star}_{\eta})\big]+\sum_{k=1}^{K}\mathbb{E}[\zeta_{k}]$ (41)
@@
$\sum_{k=2}^{K+1}\mathbb{E}[\lVert\nabla\phi_{\eta}^{\hat{\rho}}(\mathbf{x}^{k-1})\rVert^{2}]$ $=\hat{\rho}^{2}\sum_{k=2}^{K+1}\mathbb{E}[\lVert\hat{\mathbf{x}}^{k}-\mathbf{x}^{k-1}\rVert^{2}]\leq\frac{2\hat{\rho}^{2}}{\hat{\rho}-\bar{\rho}}\frac{(1+\lambda)\big[{\phi}_{\eta}(\mathbf{x}^{0})-{\phi_{\eta}}({\mathbf{x}}^{\star}_{\eta})\big]+\sum_{k=1}^{K}\mathbb{E}[\zeta_{k}]}{(1-\lambda)}.$
@@
Next, we consider solving the proximal subproblems in (33). We employ accelerated stochastic gradient descent [lan2020first, (4.2.5)-(4.2.7)]. It is a stochastic variant of the accelerated gradient descent that obtains the optimal $\mathcal{O}(L/K^{2}+\sigma^{2}/\sqrt{K})$ convergence rate for stochastic and $L$-smooth convex optimization problems. When $\sigma=0$, the algorithm reduces to the optimal gradient method with Nesterov’s acceleration.
@@
$T_{k}=\max\left\{2\sqrt{\frac{2(B+L/\eta+2\bar{\rho})}{\bar{\rho}}},\ \frac{8\sigma^{2}}{(\hat{\rho}-\bar{\rho})\varepsilon^{2}}\right\}.$
@@
$\mathcal{O}\left(\frac{\,\bar{\rho}\left[\phi(\mathbf{x}^{0})-\phi(\mathbf{x}^{\star})+R\eta\right]}{\varepsilon^{2}}\cdot\max\left\{\sqrt{\frac{B+L/\varepsilon^{2}+2\bar{\rho}}{\bar{\rho}}},\ \frac{\sigma^{2}}{\bar{\rho}\varepsilon^{2}}\right\}\right).$
@@
The construction of $T_{k}$ ensures that $\frac{4(L_{\eta}+\hat{\rho})}{\bar{\rho}T_{k}^{2}}\leq\frac{1}{2}=\lambda$ and $\zeta_{k}=\frac{4\sigma^{2}}{\bar{\rho}T_{k}}\leq\frac{\varepsilon^{2}}{2}$. Applying Theorem 5.3, we have
$\min_{1\leq k\leq K}\mathbb{E}[\lVert\nabla{\phi}_{\eta}^{\hat{\rho}}(\mathbf{x}^{k})\rVert^{2}]$ $\leq\frac{24\bar{\rho}\big[\phi(\mathbf{x}^{0})-\phi({\mathbf{x}}^{\star})+R\eta\big]}{K}+8\varepsilon^{2}.$ (44)
@@
$N=\sum_{k=1}^{K}T_{k}=\mathcal{O}\left(\frac{\bar{\rho}\big[\phi(\mathbf{x}^{0})-\phi({\mathbf{x}}^{\star})+R\eta\big]}{\varepsilon^{2}}\cdot\max\left\{\sqrt{\frac{B+L/\varepsilon^{2}+2\bar{\rho}}{\bar{\rho}}},\frac{\sigma^{2}}{\bar{\rho}\varepsilon^{2}}\right\}\right)=\mathcal{O}\Big(\max\Big\{\frac{1}{\varepsilon^{3}},\frac{\sigma^{2}}{\varepsilon^{4}}\Big\}\Big).$
@@
As an alternative to the smoothing-based approach, one may employ the proximal subgradient method [davis2019stochastic], which achieves a sample complexity of $\mathcal{O}({M^{2}}/{\varepsilon^{4}})$, where $M$ denotes the Lipschitz constant of $f$. This complexity bound is independent of the batch size. Note that in the minibatch setting (with batch size $m$), the smoothing-based method achieves a sample complexity of $\mathcal{O}(\max\{{1}/{\varepsilon^{3}},{\sigma^{2}}/(m\varepsilon^{4})\})$. It replaces the dependence on $M^{2}$ with $\sigma^{2}$ and allows minibatching to achieve variance reduction. A prior work by deng2021minibatch also considered minibatching for weakly convex optimization; however, their minibatch prox-linear and proximal point algorithms require solving increasingly complex proximal subproblems as batch size $m$ increases. In contrast, our smoothing-based approach yields simpler proximal subproblems that can be applied to each sample independently. Table 1 summarizes the resulting sample complexity guarantees.
@@
Algorithm Subgradient [davis2019stochastic] ASGD-SIPP Model-based [deng2021minibatch]
@@ 6.1 An accelerated method for generalized smooth convex optimization
where $\mathcal{L}:\mathbb{R}^{d}\times\mathbb{R}^{d}\rightarrow(0,\infty)$ is a symmetric negative continuous map.
@@
6 Early stop (optional): if $\prod_{i=1}^{t}\big(1-\sqrt{\frac{\mu}{\hat{L}_{i}+\mu}}\big)\leq\frac{1}{4}$ then return $\mathbf{y}^{t}$;
7
@@
where $\Gamma_{t}=\begin{cases}(1-\alpha_{t})^{-1}\Gamma_{t-1}&t>1\\ 1&t=1\end{cases}$, $\Delta_{t}\coloneqq\psi(\mathbf{x}^{t})-\psi(\mathbf{x}^{\star})$ and $\mathbf{x}^{\star}$ is the optimal solution.
@@
$\psi(\mathbf{y}^{T})-\psi({\mathbf{x}}^{\star})+\frac{\mu}{2}\lVert\mathbf{z}^{T}-{\mathbf{x}}^{\star}\rVert^{2}$ $\leq\prod_{1\leq t\leq T}\left(1-\sqrt{\frac{\mu}{\hat{L}_{t}+\mu}}\right)^{T}\left[\psi(\mathbf{x}^{0})-\psi({\mathbf{x}}^{\star})+\frac{\mu}{2}\lVert\mathbf{x}^{0}-{\mathbf{x}}^{\star}\rVert^{2}\right]$ (56)
@@
As it is clear from (54) that $\alpha_{t}<1$, we immediately implies $\alpha_{t}\geq\sqrt{\frac{\mu}{\hat{L}_{t}+\mu}}$. Consequently, we have
@@ 6.2 Smooth approximation algorithms for weakly convex problems
$\inf_{\mathbf{x}}\phi_{\eta}(\mathbf{x})=\inf_{\mathbf{x}}\inf_{\mathbf{y}}\left[f(\mathbf{y})+\frac{\rho+\max\{\eta^{-1},\rho\}}{2}\lVert\mathbf{y}-\mathbf{x}\rVert^{2}\right]+r(\mathbf{x})\geq\inf_{\mathbf{y}}f(\mathbf{y})+\inf_{\mathbf{x}}r(\mathbf{x})>-\infty.$
@@
Moreover, all the intermediate iterates produced by AGLS lie in the set $\mathcal{B}_{D^{\dagger}}(S_{0})\coloneqq\{\mathbf{x}:\mathrm{dist}(\mathbf{x},S_{0})\leq D^{\dagger}\},$ where $D^{\dagger}=\sqrt{\frac{8}{\hat{\rho}-\bar{\rho}}\big[{\phi}(\mathbf{x}^{0})-l\big]}$. Besides, after $K$ iterations, the total number of iterations of AGLS is bounded by $\big\lceil\log 4\sqrt{\frac{\tau_{u}\mathcal{L}^{\dagger}+\hat{\rho}-\bar{\rho}}{\hat{\rho}-\bar{\rho}}}K\big\rceil$, where $\mathcal{L}^{\dagger}=\sup\{2\mathcal{L}_{\eta}(\mathbf{x},\mathbf{y})+2\rho:\mathbf{x},\mathbf{y}\in\mathcal{B}_{{D^{\dagger}}}(S_{0})\}$.
@@
where the second equality rearranges from (58) and the last one follows from the definition of Morea envelope. Summing over $k$ yields
@@
Suppose generalized Nesterov smoothing is employed, i.e., $\phi_{\eta}(\mathbf{x})=h_{\eta}(A(\mathbf{x}))$, where $A:\mathbb{R}^{d}\!\rightarrow\!\mathbb{R}^{m}$ and each component $A_{i},i\in[m]$, is a smooth function. We assume that $h_{\eta}$ is level-bounded and that each $A_{i}$ is lower-bounded.
Remark 6.2.
The level-boundedness of $h_{\eta}$ is a natural assumption, as $h_{\eta}$ frequently serves as a loss function in machine learning. Notably, widely used examples of $h_{\eta}$, such as the softmax function and the Huber loss, possess this property.
@@
Next, for $\mathbf{x}\in\mathcal{B}_{D^{\dagger}}(S_{0})\setminus S_{0}$, we obtain by the triangle inequality
$\lVert\nabla A(\mathbf{x})\rVert_{\mathrm{op}}$ $\leq\lVert\nabla A(\mathbf{x})-\nabla A(\hat{\mathbf{x}})\rVert_{\mathrm{op}}+\lVert\nabla A(\hat{\mathbf{x}})\rVert_{\mathrm{op}}$
$\leq L_{A}\lVert\mathbf{x}-\hat{\mathbf{x}}\rVert+\lVert\nabla A(\hat{\mathbf{x}})\rVert_{\mathrm{op}}$
$\leq L_{A}D^{\dagger}+\lVert\nabla A(\hat{\mathbf{x}})\rVert_{\mathrm{op}},$

### v3

@@ 5.2 Smooth minimization with inexact proximal point
We see from Section 5.1 that directly applying smoothing does not surpass the complexity of the subgradient method [13]. The main issue is that the simple gradient descent algorithm is ineffective in dealing with the ill-conditioned smoothed problem arising from the asymmetry between the lower curvature $\bar{\rho}$ and the upper curvature $L_{\eta}$. To further improve the convergence rate, we propose solving the smooth problem using the proximal point method (Algorithm 2), which turns the original problem into a sequence of strongly convex subproblems.
@@ 2   Compute an approximate solution $\mathbf{x}^{k}$ to the following problem:
The subproblems (35) can be solved efficiently with acceleration. For convenience, we denote
@@
Let ${\mathbf{x}}^{\star}_{\eta}$ be an optimal solution to the smoothed problem. We assume that the proximal subproblem (35) is inexactly solved: the suboptimality of $\mathbf{x}^{k}$ consists of a relative error to the initial suboptimality plus an absolute error term:
@@
$\min_{1\leq k\leq K}\mathbb{E}\big[\lVert\nabla\phi_{\eta}^{\hat{\rho}}(\mathbf{x}^{k})\rVert^{2}\big]\leq\frac{2\hat{\rho}^{2}}{\hat{\rho}-\bar{\rho}}\frac{(1+\lambda)[\phi(\mathbf{x}^{0})-\phi(\mathbf{x}^{\star})+R\eta]+\sum_{k=1}^{K}\mathbb{E}[\zeta_{k}]}{(1-\lambda)K}.$
@@
$(1-\lambda)\sum_{k=2}^{K+1}\mathbb{E}\big[{\hat{\phi}}_{k}(\mathbf{x}^{k-1})-{\hat{\phi}}_{k}(\hat{\mathbf{x}}^{k})\big]$ (40)
@@
$\leq(1+\lambda)\big[{\phi_{\eta}}(\mathbf{x}^{0})-{\phi_{\eta}}({\mathbf{x}}^{\star}_{\eta})\big]+\sum_{k=1}^{K}\mathbb{E}[\zeta_{k}]$
@@
$\sum_{k=2}^{K+1}\mathbb{E}[\lVert\nabla\phi_{\eta}^{\hat{\rho}}(\mathbf{x}^{k-1})\rVert^{2}]=\hat{\rho}^{2}\sum_{k=2}^{K+1}\mathbb{E}[\lVert\hat{\mathbf{x}}^{k}-\mathbf{x}^{k-1}\rVert^{2}]\leq\frac{2\hat{\rho}^{2}}{\hat{\rho}-\bar{\rho}}\frac{(1+\lambda)[{\phi}_{\eta}(\mathbf{x}^{0})-{\phi_{\eta}}({\mathbf{x}}^{\star}_{\eta})]+\sum_{k=1}^{K}\mathbb{E}[\zeta_{k}]}{(1-\lambda)}.$
@@
Next, we consider solving the proximal subproblems in (32). We employ accelerated stochastic gradient descent [27, (4.2.5)-(4.2.7)]. It is a stochastic variant of the accelerated gradient descent that obtains the optimal $\mathcal{O}(L/K^{2}+\sigma^{2}/\sqrt{K})$ convergence rate for stochastic and $L$-smooth convex optimization problems. When $\sigma=0$, the algorithm reduces to the optimal gradient method with Nesterov’s acceleration.
@@
$T_{k}=\max\Big\{2\sqrt{\frac{2(B+L/\eta+2\bar{\rho})}{\bar{\rho}}},\ \frac{8\sigma^{2}}{\varepsilon^{2}}\Big\}.$
@@
$\mathcal{O}\Big(\frac{\,\bar{\rho}\left[\phi(\mathbf{x}^{0})-\phi(\mathbf{x}^{\star})+R\eta\right]}{\varepsilon^{2}}\cdot\max\Big\{\sqrt{\frac{B+L/\varepsilon^{2}+2\bar{\rho}}{\bar{\rho}}},\ \frac{\sigma^{2}}{\varepsilon^{2}}\Big\}\Big).$
@@
The construction of $T_{k}$ ensures that $\frac{4(L_{\eta}+\hat{\rho})}{\bar{\rho}T_{k}^{2}}\leq\frac{1}{2}=\lambda$ and $\zeta_{k}=\frac{4\sigma^{2}}{\bar{\rho}T_{k}}\leq\frac{\varepsilon^{2}}{2\bar{\rho}}$. Applying Theorem 5.3, we have
$\min_{1\leq k\leq K}\mathbb{E}[\lVert\nabla{\phi}_{\eta}^{\hat{\rho}}(\mathbf{x}^{k})\rVert^{2}]$ $\leq\frac{24\bar{\rho}[\phi(\mathbf{x}^{0})-\phi({\mathbf{x}}^{\star})+R\eta]}{K}+8\varepsilon^{2}.$ (43)
@@
$N=\sum_{k=1}^{K}T_{k}=\mathcal{O}\left(\frac{\bar{\rho}\big[\phi(\mathbf{x}^{0})-\phi({\mathbf{x}}^{\star})+R\eta\big]}{\varepsilon^{2}}\cdot\max\left\{\sqrt{\frac{B+L/\varepsilon^{2}+2\bar{\rho}}{\bar{\rho}}},\frac{\sigma^{2}}{\varepsilon^{2}}\right\}\right)=\mathcal{O}\Big(\max\Big\{\frac{1}{\varepsilon^{3}},\frac{\sigma^{2}}{\varepsilon^{4}}\Big\}\Big).$
@@
As an alternative to the smoothing-based approach, one may employ the proximal subgradient method [13], which achieves a sample complexity of $\mathcal{O}({M^{2}}/{\varepsilon^{4}})$, where $M$ denotes the Lipschitz constant of $f$. This complexity bound is independent of the batch size. Note that in the minibatch setting (with batch size $m$), the smoothing-based method achieves a sample complexity of $\mathcal{O}(\max\{{1}/{\varepsilon^{3}},{\sigma^{2}}/(m\varepsilon^{4})\})$. It replaces the dependence on $M^{2}$ with $\sigma^{2}$ and allows minibatching to achieve variance reduction. A prior work by Deng and Gao [15] also considered minibatching for weakly convex optimization; however, their minibatch prox-linear and proximal point algorithms require solving increasingly complex proximal subproblems as batch size $m$ increases. In contrast, our smoothing-based approach yields simpler proximal subproblems that can be applied to each sample independently. Table 1 summarizes the resulting sample complexity guarantees.
@@
Algorithm Subgradient [13] ASGD-SIPP Model-based [15]
@@
Remark 5.3.
Instead of using inexact proximal point, it is also feasible to adopt other algorithms that can exploit the asymmetry of the function’s upper/lower curvature. For example, adopting the parameter-free algorithm from [26] yields an $\mathcal{O}\big(\frac{\sqrt{L_{\eta}\bar{\rho}}}{\varepsilon^{2}}\big)=\mathcal{O}(\frac{1}{\varepsilon^{3}})$ complexity without knowing $\bar{\rho}$.
@@ 6.1 An accelerated method for generalized smooth convex optimization
where $\mathcal{L}:\mathbb{R}^{d}\times\mathbb{R}^{d}\rightarrow(0,\infty)$ is a symmetric non-negative continuous map.
@@
Early stop (optional): if $\prod_{i=1}^{t}\big(1-\sqrt{\frac{\mu}{\hat{L}_{i}+\mu}}\big)\leq\frac{1}{4}$ then return $\mathbf{y}^{t}$;
6
@@
where $\Gamma_{t}=\begin{cases}(1-\alpha_{t})^{-1}\Gamma_{t-1}&t>1\\ 1&t=1\end{cases}$, $\Delta_{t}\coloneqq\psi(\mathbf{y}^{t})-\psi(\mathbf{x}^{\star})$ and $\mathbf{x}^{\star}$ is the optimal solution.
@@
$\psi(\mathbf{y}^{T})-\psi({\mathbf{x}}^{\star})+\frac{\mu}{2}\lVert\mathbf{z}^{T}-{\mathbf{x}}^{\star}\rVert^{2}$ $\leq\prod_{1\leq t\leq T}\left(1-\sqrt{\frac{\mu}{\hat{L}_{t}+\mu}}\right)\left[\psi(\mathbf{x}^{0})-\psi({\mathbf{x}}^{\star})+\frac{\mu}{2}\lVert\mathbf{x}^{0}-{\mathbf{x}}^{\star}\rVert^{2}\right]$ (55)
@@
As it is clear from (53) that $\alpha_{t}<1$, we immediately have $\alpha_{t}\geq\sqrt{\frac{\mu}{\hat{L}_{t}+\mu}}$. Consequently, we have
@@ 6.2 Smooth approximation algorithms for weakly convex problems
$\inf_{\mathbf{x}}\phi_{\eta}(\mathbf{x})=\inf_{\mathbf{x}}\inf_{\mathbf{y}}\big[f(\mathbf{y})+\frac{\rho+\max\{\eta^{-1},\rho\}}{2}\lVert\mathbf{y}-\mathbf{x}\rVert^{2}\big]+r(\mathbf{x})\geq\inf_{\mathbf{y}}f(\mathbf{y})+\inf_{\mathbf{x}}r(\mathbf{x})>-\infty.$
@@
Moreover, all the intermediate iterates produced by AGLS lie in the set $\mathcal{B}_{D^{\dagger}}(S_{0})\coloneqq\{\mathbf{x}:\mathrm{dist}(\mathbf{x},S_{0})\leq D^{\dagger}\},$ where $D^{\dagger}=\sqrt{\frac{8}{\hat{\rho}-\bar{\rho}}\big[{\phi}(\mathbf{x}^{0})-l\big]}$. Besides, after $K$ iterations, the total number of iterations of AGLS is bounded by $\big\lceil\log 4\sqrt{\frac{\tau_{u}\mathcal{L}^{\dagger}+\hat{\rho}-\bar{\rho}}{\hat{\rho}-\bar{\rho}}}K\big\rceil$, where $\mathcal{L}^{\dagger}=\sup\{2\mathcal{L}_{\eta}(\mathbf{x},\mathbf{y})+2\bar{\rho}:\mathbf{x},\mathbf{y}\in\mathcal{B}_{{D^{\dagger}}}(S_{0})\}$.
@@
where the second equality rearranges from (57) and the last one follows from the definition of $\hat{\phi}_{k}$ and nonnegativity of the quadratic term. Summing over $k$ yields
@@
Suppose generalized Nesterov smoothing is employed, i.e., $\phi_{\eta}(\mathbf{x})=f_{|}eta(\mathbf{x})=h_{\eta}(A(\mathbf{x}))$, where $A:\mathbb{R}^{d}\!\rightarrow\!\mathbb{R}^{m}$ and each component $A_{i},i\in[m]$, is a smooth function. We assume that $h_{\eta}$ is level-bounded and that each $A_{i}$ is lower-bounded.
The level-boundedness of $h_{\eta}$ is a natural assumption, as $h_{\eta}$ frequently serves as a loss function in machine learning. Notably, widely used examples of $h_{\eta}$, such as the Huber loss, possess this property.
@@
Next, for $\mathbf{x}\in\mathcal{B}_{D^{\dagger}}(S_{0})\setminus S_{0}$, we obtain by the triangle inequality that
$\lVert\nabla A(\mathbf{x})\rVert_{\mathrm{op}}$ $\leq\lVert\nabla A(\mathbf{x})-\nabla A(\hat{\mathbf{x}})\rVert_{\mathrm{op}}+\lVert\nabla A(\hat{\mathbf{x}})\rVert_{\mathrm{op}}\leq L_{A}\lVert\mathbf{x}-\hat{\mathbf{x}}\rVert+\lVert\nabla A(\hat{\mathbf{x}})\rVert_{\mathrm{op}}\leq L_{A}D^{\dagger}+\lVert\nabla A(\hat{\mathbf{x}})\rVert_{\mathrm{op}},$